OpenStudy (anonymous):

sum of two numbers is 6 times their geometric means , show that the numbers are in this ratio

OpenStudy (anonymous):

ratio is \[3+2\sqrt{2}:3- 2\sqrt{2}\]

OpenStudy (anonymous):

let gm be a and the numbers be x and y a^2 = xy x + y = 6xy

OpenStudy (anonymous):

sorry x + y = 6(xy)^1/2

OpenStudy (anonymous):

dont you have any other information

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x + y - 6(xy)^1/2 = 0 (x^1/2 - y^1/2)^2 = 4(xy)^1/2

OpenStudy (anonymous):

rhs = x -x + y - y + 2(xy)^1/2 + 2 (xy)^1/2

OpenStudy (anonymous):

rhs = -(x^1/2 - y^1/2)^2 + (x^1/2 + y^1/2)^2

OpenStudy (anonymous):

2 ( x^1/2 - y^1/2)^2 = (x^1/2 + y^1/2)^2

OpenStudy (anonymous):

a + b = 6*\[\sqrt{ab}\] squaring... a^2 + b^2 +2ab=36ab ..divide by b^2 (a/b)^2 -34(a/b) +1=0.. solving quadratic eqn a/b=33.970 reqd...ratio

OpenStudy (anonymous):

x^1/2 + y^1/2 2^1/2 = -------------------------- x^1/2 - y^1/2

OpenStudy (anonymous):

2^1/2 + 1 2 x^1/2 --------- = ------------ 2^1/2 - 1 2 y^1/2 square both sides 2.2^1/2 + 3 x ------------- = ----- 3 -2.2^1/2 y

OpenStudy (anonymous):

there goes your answer : )

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

ok